# ACDC: Automated Creation of Digital Cousins for Robust Policy Learning

### v2

@@ Abstract
Training robot policies in the real world can be unsafe, costly, and difficult to scale. Simulation serves as an inexpensive and potentially limitless source of training data, but suffers from the semantics and physics disparity between simulated and real-world environments. These discrepancies can be minimized by training in digital twins, which serve as virtual replicas of a real scene but are expensive to generate and cannot produce cross-domain generalization. To address these limitations, we propose the concept of digital cousins, a virtual asset or scene that, unlike a digital twin, does not explicitly model a real-world counterpart but still exhibits similar geometric and semantic affordances. As a result, digital cousins simultaneously reduce the cost of generating an analogous virtual environment while also facilitating better robustness during sim-to-real domain transfer by providing a distribution of similar training scenes. Leveraging digital cousins, we introduce a novel method for the Automatic Creation of Digital Cousins (ACDC), and propose a fully automated real-to-sim-to-real pipeline for generating fully interactive scenes and training robot policies that can be deployed zero-shot in the original scene. We find that ACDC can produce digital cousin scenes that preserve geometric and semantic affordances, and can be used to train policies that outperform policies trained on digital twins, achieving 90% vs. 25% under zero-shot sim-to-real transfer. Additional details are available at https://digital-cousins.github.io/.
@@
Figure 1: Overview. ACDC is a completely automatic pipeline generating fully interactive simulated scenes (what we call digital cousins) from a single RGB image. Unlike a digital twin, digital cousins relax the assumption of completely reconstructing the minute details of a given scene and instead focus on preserving higher-level details, such as spatial relationships and semantic affordances. By leveraging motion planning and ground-truth simulation information, we can automatically collect demonstrations in our digital cousin scenes, augmented with physically plausible randomizations. A policy trained on these synthetic demonstrations can then be deployed zero-shot in the original scene, without requiring any additional finetuning.
@@ 1 Introduction
Leveraging digital cousins, we then introduce a novel method for the Automatic Creation of Digital Cousins (ACDC) that can be used fully-automated end-to-end in a real-to-sim-to-real setup, in which digital cousins generated from a real-world image can be used to train policies that can be deployed zero-shot in the original scene. ACDC leverages DINOv2 [11] as a proxy for measuring similarities between a given real-world asset and candidate digital assets, as it has been shown to visually encode relevant geometric and spatial information from diverse sets of images, and we consider assets with low feature embedding distances as being digital cousins of a given real-world object.
Our contributions are threefold. First, we propose ACDC, a novel method requiring zero human input to generate digital cousin scenes from a single image. Second, we provide an automated recipe to train simulation policies in digital cousins generated by ACDC. Third, we show that robot manipulation policies trained within digital cousins can match the performance of those trained on digital twins, and can outperform digital twin policies when tested on unseen objects, both in sim and in the original real-world scene. Code and videos can be found in supplementary materials and on the project website https://digital-cousins.github.io/.
@@ 2 Methodology
2.1 Automatic Creation of Digital Cousins (ACDC)
@@
We additionally require a depth map in order to properly position and rescale matched digital cousins when generating our scene. Depth cameras are widely used but cannot accurately capture reflective surfaces and prevent ACDC from being used on in-the-wild images. To mitigate these limitations, we leverage Depth-Anything-v2 [14], a state-of-the-art monocular depth estimation model, to estimate the corresponding depth map $\mathbf{D}$ from $\mathbf{X}$. We then extract point cloud $\mathbf{P}=\mathbf{D}\cdot\mathbf{K}^{-1}$, and leverage individual object masks $\mathbf{m}_{i}$ to generate the subset of points $\mathbf{p_{i}}$ from $\mathbf{P}$ and pixels $\mathbf{x_{i}}$ from $\mathbf{X}$ corresponding to that object, resulting in a set of object representations $\{\mathbf{o}_{i}=(\mathbf{l}_{i},\mathbf{m}_{i},\mathbf{p}_{i},\mathbf{x}_{i})\}_{i=1}^{N}$.
@@
Our final step in ACDC is to compile our matched cousins into a physically plausible digital cousin scene. For given input object information $\mathbf{o}_{i}$ and corresponding matched cousin information $(\mathbf{A}_{c},\mathbf{q}_{c})$, we place the asset’s bounding box center at the centroid of the corresponding input object point cloud $\mathbf{p}_{i}$, and then rescale to align with $\mathbf{p}_{i}$’s extents. We additionally fit floor and wall planes from their obtained point clouds from the extraction step, and query GPT-4 to determine whether any objects should be mounted on either the floor or wall. Finally, we de-penetrate all objects so that the scene is physically stable. For additional scene post-processing details, please see Section A.3.
@@ 2.2 Policy Learning
To facilitate automatic demonstration collection in simulation, we implement a set of sample-based skills that leverage both motion planning and ground-truth simulation data. Concretely, our skills include Open, Close, Pick, and Place. For specific implementation details, please see Section A.4. While currently limited, these skills already enable demonstration collection across a wide range of everyday tasks, such as object rearrangement and furniture articulation.
@@
Input Scene ACDC Output Scale (m) Cat. Mod. $\mathcal{L}_{2}$ Dist. (cm) $\downarrow$ Ori. Diff. (rad) $\downarrow$ Bbox IoU $\uparrow$ Cen. IoU $\uparrow$
@@
Table 1: Quantitative and qualitative evaluation of scene reconstruction by ACDC in a sim-to-sim scenario. ‘Scale’ is the largest distance between two objects in the input scene. ‘Cat.’ indicates the ratio of correctly categorized objects to the total number of objects in the scene. ‘Mod.’ shows the ratio of correctly modeled objects to the total number of objects. ‘$\mathcal{L}_{2}$ Dist.’ provides the mean and standard deviation of the Euclidean distance between the centers of the bounding boxes in the input and reconstructed scenes. ‘Ori. Diff.’ represents the mean and standard deviation of the orientation magnitude difference of each centrosymmetric object. ‘Bbox IoU’ presents the Intersection over Union (IoU) for assets’ 3D bounding boxes. ‘Cen. IoU’ shows the IoU for assets’ 3D bounding boxes after aligning their center position. Please refer to Section B.1 for more results.
@@ 3 Experiments
  1. Q1. Can ACDC produce high-quality digital cousin scenes? Given a single RGB image, can ACDC capture the high-level semantic and spatial details inherent in the original scene?
@@
3.1 Scene Reconstruction via ACDC
@@
We will show quantitative evaluation and qualitative results of ACDC to answer Q1. To quantify the reconstructive ability of ACDC, we first test our method on a variety of simulated and real scenes to perform both sim-to-sim and real-to-sim scene reconstruction, where we input a single RGB image of a simulated scene and generate the closest digital cousins using ACDC. In the sim-to-sim setup, we have guaranteed access to its “digital twin” (i.e., the ground truth category and model), as well as ground truth information about all scene objects’ poses and scales, and can quantitatively measure ACDC’s reconstructive fidelity. In this setting, we measure the proportion of scene objects whose category and model were successfully preserved in the nearest digital cousin to capture the digital cousin’s semantic fidelity, and measure the averaged per-object pose error (via $\mathcal{L}_{2}$ distance and orientation difference) and scale error (via bounding box IoU) to capture its geometric fidelity. However, we do not have access to digital twins for the real-world objects nor ground truth spatial information; instead, we provide qualitative side-by-side comparisons between the real-world scene and its corresponding digital cousin scene. Sim-to-sim results in Table 1, real-to-sim results in Fig. 3, and additional results in Appendix B.
Figure 3: Qualitative ACDC real-to-sim scene reconstruction results. Multiple cousins are shown with a robot collecting demonstrations. Please refer to Section B.2 for more results.
Digital cousin generation: ACDC can preserve semantic and spatial details (Q1). In the sim-to-sim setup, we find that ACDC correctly reproduces the original per-object category and model in most cases. Spatially, we also find that ACDC can scale and place objects that similarly match their original counterparts in the input scene. Qualitatively, the side-by-side comparison of our input- and ACDC-generated scenes showcase the immediate visual similarity between the two, and suggest that our quantitative results imply a digital cousin scene quality that can successfully preserve the original scene’s object layout. In the real-to-sim setup, find that ACDC produces reasonable scenes that are both physically plausible and able to preserve scene-level semantic and spatial details.
@@
Based on these results, we can safely answer Q1: ACDC can indeed preserve semantic and spatial details of input scenes, generating cousins of real-world objects from a single RGB image that can be accurately positioned and scaled to match the original scene.
@@ 3.4 Real-to-Sim-to-Real Scene Generation and Policy Learning
Finally, we test our full ACDC pipeline and automated policy learning framework end-to-end with a fully in-the-wild kitchen scene. We find that our policy can successfully open the kitchen cabinet after being trained exclusively in simulation on digital cousins, as seen in Fig. 1. Experiment videos and additional results can be found at https://digital-cousins.github.io/.
@@ 4 Related Work
Creating realistic and diverse digital assets and scenes from real-world inputs is a prevalent and long-standing problem [16, 17, 18, 19]. Within robot learning, real-to-sim scene creation has been achieved through manual curation [20, 21, 22, 23, 24, 25, 4, 10], procedural generation [26, 27, 28], few-shot interactions [29, 8, 30], inverse graphics [31], and more recently foundation model-assisted generation [32, 33]. However, these methods either cannot handle scene-level generation, require human labor, or cannot retain physical plausibility. In contrast, ACDC is fully automated and is faithful to the input physical scenes.
@@ 5 Conclusion
ACDC is a fully automated pipeline used to quickly generate fully interactive digital cousin scenes corresponding to a single real-world RGB image. We find that policies trained on these digital cousin setups are more robust than those trained on digital twins, with comparable in-domain performance and superior out-of-domain generalization, and enable zero-shot sim-to-real policy transfer.
Limitations. Our system has a few limitations. First, ACDC is bounded by the diversity of its underlying asset dataset. While BEHAVIOR-1K contains thousands of unique assets, we find that it is still insufficient to densely capture the real-world distribution of objects. Second, because ACDC is built upon multiple large pretrained models, our pipeline inherits the limitations of these models, including adversarial and out-of-distribution scenes. Third, policy learning with ACDC can still be significantly improved and can benefit from recent advancements in robot learning, such as diffusion policies [85].
@@ Acknowledgments
We are grateful to the SVL PAIR group for helpful feedback and insightful discussions. This work is in part supported by the Stanford Institute for Human-Centered AI (HAI), ONR MURI N00014-22-1-2740, ONR MURI N00014-21-1-2801, and Schmidt Sciences. Ruohan Zhang is partially supported by the Wu Tsai Human Performance Alliance Fellowship.
@@
Appendix A Additional ACDC Details
@@ A.7 Additional Real-to-Sim Details
One key design decision in ACDC is our decision to use synthetic depth via Depth-Anything-v2 [14], instead of a dedicated depth camera. This decision is guided by our observation that it performs more consistently on reflective surfaces. However, this synthetic depth approach still generates artifacts occurring near object boundaries and the image periphery, and so to mitigate this issue we post-process the output of Depth-Anything-v2 by applying an erosion kernel to segmented object masks $\mathbf{m}_{i}$. To further remove noise in object point clouds, we apply DBSCAN clustering [93] on each object point cloud $\mathbf{p}_{i}$ to filter out noisy points.
@@
DINO performs a rough estimation of asset orientations, which for most objects the orientation is sufficiently accurate. However, ACDC additionally provides an option to further refine the orientation refinement based on an object’s extracted point cloud. By computing the z-aligned minimum bounding box of the given point cloud, we can apply an additional z-rotation to DINO’s outputted estimated orientation so that the matched asset’s canonical xy-axes aligns with the computed minimum bounding box frame. We find this is especially useful for object’s that have sharp geometric boundaries, such as furniture objects.
@@ B.1 Visual Encoder Ablation Study
Input Scene ACDC Output Scale (m) Cat. Mod. L2 Dist (cm) $\downarrow$ Ori. Diff. $\downarrow$ Bbox IoU $\uparrow$ Ori. Bbox IoU $\uparrow$
@@
Table 2: Quantitative evaluation of scene reconstruction using our ACDC method in a sim-to-sim scenario. This table is an extension of Table 1 in the main paper. ‘Cat.’ indicates the ratio of correctly categorized objects to the total number of objects in the scene. ‘Mod.’ shows the ratio of correctly modeled objects to the total number of objects in the scene. ‘L2 Dist’ provides the mean and standard deviation of the Euclidean distance between the centers of the bounding boxes in the input and reconstructed scenes. ‘Ori. Diff.’ represents the mean and standard deviation of the orientation magnitude difference of each non-uniformly symmetric object. ‘Bbox IoU’ presents the Intersection over Union (IoU) for axis-aligned 3D bounding boxes. ‘Ori. Bbox IoU’ displays the IoU for oriented 3D bounding boxes.
Figure 6: Qualitative ACDC sim-to-sim scene reconstruction results. Overall, pipeline (d) gives the best scene reconstruction results, while pipeline (c) balances inference time and reconstruction quality.
@@
Table 2 presents a quantitative evaluation of our ACDC in the sim-to-sim setting, while Fig. 6 provides qualitative visualizations of the output scenes for each pipeline. To ensure diversity at the object level, no model is present in more than one test scene.
@@
When comparing (b) with (d), the differences in orientation difference and IoU metrics are minimal, indicating that ACDC can reconstruct high-quality scenes even when the assets in the simulated scene are close approximations (cousins) rather than exact replicas (twins) of the target objects.
Finally, examining the L2 Dist column in Table 2, we see that ACDC places each asset very close to the ground truth position. The average L2 distance errors are less than 10 cm for the first seven test scenes, and is only 17 cm for the eighth scene whose scale is 10.23 m.

### v3

# Automated Creation of Digital Cousins for Robust Policy Learning

@@ Abstract
Training robot policies in the real world can be unsafe, costly, and difficult to scale. Simulation serves as an inexpensive and potentially limitless source of training data, but suffers from the semantics and physics disparity between simulated and real-world environments. These discrepancies can be minimized by training in digital twins, which serve as virtual replicas of a real scene but are expensive to generate and cannot produce cross-domain generalization. To address these limitations, we propose the concept of digital cousins, a virtual asset or scene that, unlike a digital twin, does not explicitly model a real-world counterpart but still exhibits similar geometric and semantic affordances. As a result, digital cousins simultaneously reduce the cost of generating an analogous virtual environment while also facilitating better robustness during sim-to-real domain transfer by providing a distribution of similar training scenes. Leveraging digital cousins, we introduce a novel method for their automated creation, and propose a fully automated real-to-sim-to-real pipeline for generating fully interactive scenes and training robot policies that can be deployed zero-shot in the original scene. We find that digital cousin scenes that preserve geometric and semantic affordances can be produced automatically, and can be used to train policies that outperform policies trained on digital twins, achieving $90\%$ vs. $25\%$ success rates under zero-shot sim-to-real transfer. Additional details are available at https://digital-cousins.github.io/.
@@
Figure 1: Overview. Fully interactive digital cousin scenes can be generated completely automatically from a single RGB image. Unlike a digital twin, digital cousins relax the assumption of completely reconstructing the minute details of a given scene and instead focus on preserving higher-level details, such as spatial relationships and semantic affordances. By leveraging motion planning and ground-truth simulation information, we can automatically collect demonstrations in our digital cousin scenes, augmented with physically plausible randomizations. A policy trained on these synthetic demonstrations can then be deployed zero-shot in the original scene, without requiring any additional finetuning.
@@ 1 Introduction
Leveraging digital cousins, we then introduce a novel method for the Automated Creation of Digital Cousins (ACDC) that can be used fully-automated end-to-end in a real-to-sim-to-real setup, in which digital cousins generated from a real-world image can be used to train policies deployed zero-shot in the original scene. ACDC leverages DINOv2 [11] as a proxy for measuring similarities between a given real-world asset and candidate digital assets, as it has been shown to visually encode relevant geometric and spatial information from diverse sets of images, and we consider assets with low feature embedding distances as being digital cousins of a given real-world object.
Our contributions are threefold. First, we propose the concept of digital cousins and a novel method ACDC for their automated creation from a single image requiring zero human input. Second, we provide an automated recipe to train simulation policies in digital cousins. Third, we show that robot manipulation policies trained within digital cousins can match the performance of those trained on digital twins, and can outperform digital twin policies when tested on unseen objects, both in simulation and in the original real-world scene. Code and videos can be found on the project website https://digital-cousins.github.io/.
@@ 2 Methodology
2.1 Automated Creation of Digital Cousins (ACDC)
@@
We additionally require a depth map in order to properly position and rescale matched digital cousins when generating our scene. Depth cameras are widely used but cannot accurately capture reflective surfaces and prevent usage on in-the-wild images. To mitigate these limitations, we leverage Depth-Anything-v2 [14], a state-of-the-art monocular depth estimation model, to estimate the corresponding depth map $\mathbf{D}$ from $\mathbf{X}$. We then extract point cloud $\mathbf{P}=\mathbf{D}\cdot\mathbf{K}^{-1}$, and leverage individual object masks $\mathbf{m}_{i}$ to generate the subset of points $\mathbf{p_{i}}$ from $\mathbf{P}$ and pixels $\mathbf{x_{i}}$ from $\mathbf{X}$ corresponding to that object, resulting in a set of object representations $\{\mathbf{o}_{i}=(\mathbf{l}_{i},\mathbf{m}_{i},\mathbf{p}_{i},\mathbf{x}_{i})\}_{i=1}^{N}$.
@@
The final step is to compile our matched cousins into a physically plausible digital cousin scene. For given input object information $\mathbf{o}_{i}$ and corresponding matched cousin information $(\mathbf{A}_{c},\mathbf{q}_{c})$, we place the asset’s bounding box center at the centroid of the corresponding input object point cloud $\mathbf{p}_{i}$, and then rescale to align with $\mathbf{p}_{i}$’s extents. We additionally fit floor and wall planes from their obtained point clouds from the extraction step, and query GPT-4 to determine whether any objects should be mounted on either the floor or wall. Finally, we de-penetrate all objects so that the scene is physically stable. For additional scene post-processing details, please see Section A.3.
@@ 2.2 Policy Learning
To facilitate automated demonstration collection in simulation, we implement a set of sample-based skills that leverage both motion planning and ground-truth simulation data. Concretely, our skills include Open, Close, Pick, and Place. For specific implementation details, please see Section A.4. While currently limited, these skills already enable demonstration collection across a wide range of everyday tasks, such as object rearrangement and furniture articulation.
@@
Input Scene Cousin Scene Scale (m) Cat. Mod. $\mathcal{L}_{2}$ Dist. (cm) $\downarrow$ Ori. Diff. (rad) $\downarrow$ Bbox IoU $\uparrow$ Cen. IoU $\uparrow$
@@
Table 1: Quantitative and qualitative evaluation of nearest digital cousin scene reconstruction in a sim-to-sim scenario. ‘Scale’ is the largest distance between two objects in the input scene. ‘Cat.’ indicates the ratio of correctly categorized objects to the total number of objects in the scene. ‘Mod.’ shows the ratio of correctly modeled objects to the total number of objects. ‘$\mathcal{L}_{2}$ Dist.’ provides the mean and standard deviation of the Euclidean distance between the centers of the bounding boxes in the input and reconstructed scenes. ‘Ori. Diff.’ represents the mean and standard deviation of the orientation magnitude difference of each centrosymmetric object. ‘Bbox IoU’ presents the Intersection over Union (IoU) for assets’ 3D bounding boxes. ‘Cen. IoU’ shows the IoU for assets’ 3D bounding boxes after aligning their center position. Please refer to Section B.1 for more results.
@@ 3 Experiments
  1. Q1. Can ACDC produce high-quality digital cousin scenes? Given a single RGB image, can the recovered digital cousins capture the high-level semantic and spatial details inherent in the original scene?
@@
3.1 Digital Cousin Scene Generation via ACDC
@@
We will show quantitative evaluation and qualitative results of recovered digital cousins to answer Q1. To quantify the quality of generated digital cousins, we first test our method on a variety of simulated and real scenes to perform both sim-to-sim and real-to-sim digital cousin scene generation, where we input a single RGB image of a simulated scene and generate the closest digital cousins using ACDC. In the sim-to-sim setup, we have guaranteed access to its “digital twin” (i.e., the ground truth category and model), as well as ground truth information about all scene objects’ poses and scales, and can quantitatively measure the reconstructive fidelity. In this setting, we measure the proportion of scene objects whose category and model were successfully preserved in the nearest digital cousin to capture the digital cousin’s semantic fidelity, and measure the averaged per-object pose error (via $\mathcal{L}_{2}$ distance and orientation difference) and scale error (via bounding box IoU) to capture its geometric fidelity. However, we do not have access to digital twins for the real-world objects nor ground truth spatial information; instead, we provide qualitative side-by-side comparisons between the real-world scene and its corresponding digital cousin scenes. Sim-to-sim results in Table 1, real-to-sim results in Fig. 3, and additional results in Appendix B.
Figure 3: Qualitative real-to-sim digital cousin scene generation results. Multiple cousins are shown with a robot collecting demonstrations. Please refer to Section B.2 for more results.
Digital cousin generation: Semantic and spatial details are preserved (Q1). In the sim-to-sim setup, we find that the original per-object category and model are correctly reproduced in most cases. Spatially, we also find that scales and positions of reconstructed digital cousins can similarly match their original counterparts in the input scene. Qualitatively, the side-by-side comparison of our input- and ACDC-generated scenes showcase the immediate visual similarity between the two, and suggest that our quantitative results imply a digital cousin scene quality that can successfully preserve the original scene’s object layout. In the real-to-sim setup, we find that ACDC produces reasonable scenes that are both physically plausible and able to preserve scene-level semantic and spatial details.
@@
Based on these results, we can safely answer Q1: digital cousins can indeed preserve semantic and spatial details of input scenes, reconstructed from a single RGB image that can be accurately positioned and scaled to match the original scene.
@@ 3.4 Real-to-Sim-to-Real Scene Generation and Policy Learning
Finally, we test our full pipeline and automated policy learning framework end-to-end with a fully in-the-wild kitchen scene. We find that our policy can successfully open the kitchen cabinet after being trained exclusively in simulation on digital cousins, as seen in Fig. 1. Experiment videos and additional results can be found at https://digital-cousins.github.io/.
@@ 4 Related Work
Creating realistic and diverse digital assets and scenes from real-world inputs is a prevalent and long-standing problem [16, 17, 18, 19]. Within robot learning, real-to-sim scene creation has been achieved through manual curation [20, 21, 22, 23, 24, 25, 4, 10], procedural generation [26, 27, 28], few-shot interactions [29, 8, 30], inverse graphics [31], and more recently foundation model-assisted generation [32, 33]. However, these methods either cannot handle scene-level generation, require human labor, or cannot retain physical plausibility. In contrast, ACDC is fully automated and the recovered digital cousins are faithful to the input physical scenes.
@@ 5 Conclusion
Digital cousins can be quickly generated by a fully automated pipeline, called ACDC, from a single real-world RGB image. We find that policies trained on these digital cousins are more robust than those trained on digital twins, with comparable in-domain performance and superior out-of-domain generalization, and enable zero-shot sim-to-real policy transfer.
Limitations. Our system has a few limitations. First, ACDC is bounded by the diversity of its underlying asset dataset. While BEHAVIOR-1K contains thousands of unique assets, we find that it is still insufficient to densely capture the real-world distribution of objects. Second, because ACDC is built upon multiple large pretrained models, our pipeline inherits the limitations of these models, including adversarial and out-of-distribution scenes. Third, policy learning with digital cousins can still be significantly improved and can benefit from recent advancements in robot learning, such as diffusion policies [85].
@@ Acknowledgments
We are grateful to the SVL PAIR group for helpful feedback and insightful discussions. This work is in part supported by the Stanford Institute for Human-Centered AI (HAI), ONR MURI N00014-22-1-2740, ONR YIP N00014-24-1-2117, ONR MURI N00014-21-1-2801, and Schmidt Sciences. Ruohan Zhang is partially supported by the Wu Tsai Human Performance Alliance Fellowship.
@@
Appendix A Additional Cousin Creation Details
@@ A.7 Additional Real-to-Sim Details
One key design decision is to use synthetic depth via Depth-Anything-v2 [14], instead of a dedicated depth camera. This decision is guided by our observation that it performs more consistently on reflective surfaces. However, this synthetic depth approach still generates artifacts occurring near object boundaries, the image periphery, and under lighting changes. To further remove noise in object point clouds, we apply DBSCAN clustering [93] on each object point cloud $\mathbf{p}_{i}$ to filter out noisy points.
@@
DINO performs a rough estimation of asset orientations, which for most objects the orientation is sufficiently accurate. However, we additionally provide an option to further refine the orientation refinement based on an object’s extracted point cloud. By computing the z-aligned minimum bounding box of the given point cloud, we can apply an additional z-rotation to DINO’s outputted estimated orientation so that the matched asset’s canonical xy-axes aligns with the computed minimum bounding box frame. We find this is especially useful for object’s that have sharp geometric boundaries, such as furniture objects.
@@ B.1 Visual Encoder Ablation Study
Input Scene Cousin Scene Scale (m) Cat. Mod. L2 Dist (cm) $\downarrow$ Ori. Diff. $\downarrow$ Bbox IoU $\uparrow$ Ori. Bbox IoU $\uparrow$
@@
Table 2: Quantitative evaluation of nearest digital cousin scene reconstruction in a sim-to-sim scenario. This table is an extension of Table 1 in the main paper. ‘Cat.’ indicates the ratio of correctly categorized objects to the total number of objects in the scene. ‘Mod.’ shows the ratio of correctly modeled objects to the total number of objects in the scene. ‘L2 Dist’ provides the mean and standard deviation of the Euclidean distance between the centers of the bounding boxes in the input and reconstructed scenes. ‘Ori. Diff.’ represents the mean and standard deviation of the orientation magnitude difference of each non-uniformly symmetric object. ‘Bbox IoU’ presents the Intersection over Union (IoU) for axis-aligned 3D bounding boxes. ‘Ori. Bbox IoU’ displays the IoU for oriented 3D bounding boxes.
Figure 6: Qualitative sim-to-sim digital cousin scene reconstruction results. Overall, pipeline (d) gives the best scene reconstruction results, while pipeline (c) balances inference time and reconstruction quality.
@@
Table 2 presents a quantitative evaluation of our digital cousin creation in the sim-to-sim setting, while Fig. 6 provides qualitative visualizations of the output scenes for each pipeline. To ensure diversity at the object level, no model is present in more than one test scene.
@@
When comparing (b) with (d), the differences in orientation difference and IoU metrics are minimal, indicating that high-quality scenes can be reconstructed even when the assets in the simulated scene are close approximations (cousins) rather than exact replicas (twins) of the target objects.
Finally, examining the L2 Dist column in Table 2, we see that each asset is placed very close to the ground truth position. The average L2 distance errors are less than 10 cm for the first seven test scenes, and is only 17 cm for the eighth scene whose scale is 10.23 m.